OpenStudy (anonymous):

f(x) = x - 4 g(x) = 3x + 6 f(5) + g(1) / f(6) - g(-2) ??

OpenStudy (anonymous):

What is f(5)?

OpenStudy (anonymous):

I don't know.

OpenStudy (anonymous):

f(5) is the value of the function f, when x = 5. So you tell me, what is f(5)?

OpenStudy (anonymous):

well?

OpenStudy (anonymous):

so f(5) =5-4

OpenStudy (anonymous):

great..which is 1

OpenStudy (anonymous):

what is g(1)?

OpenStudy (anonymous):

3(1) + 6 = 9

OpenStudy (anonymous):

Yes! So now you can do the question you asked. Do it and I will tell you if you are right or wrong.

OpenStudy (anonymous):

I got 10/2 = 5 as my final answer

OpenStudy (anonymous):

Correct! Way to go!